OpenStudy (kc_kennylau):

Express this function using defined step-functions please? \[r=\left\{\begin{array}{lr} |\sec\theta| &\mbox{if}&-\frac\pi4<\theta<\frac\pi4 &\mbox{or}&\frac{3\pi}4<\theta<\frac{5\pi}4\\ |\csc\theta| &\mbox{if}&\frac\pi4<\theta<\frac{3\pi}4 &\mbox{or}&\frac{5\pi}4<\theta<\frac{7\pi}4 \end{array}\right.\]

OpenStudy (kc_kennylau):

@primeralph how to disable and enable functions? :O

OpenStudy (anonymous):

The piece-wise function\[ f(t) = \begin{cases} g(t)& t>a\\ h(t) &t<a \end{cases} \]Can be expressed as:\[ f(t) =h(t)+[g(t)-h(t)]H(t-a) \] Where \(H(t)\) is the Heaviside function:\[ H(t) = \begin{cases} 1& t>0\\ 0& t<0 \end{cases} \]

OpenStudy (anonymous):

Just keep expanding out using this equation.

OpenStudy (kc_kennylau):

Is there no other way than using Heaviside? :/

OpenStudy (anonymous):

hey @kc_kennylau!! long time no see it's been about 5 hours!

OpenStudy (kc_kennylau):

@DLover243 hi :)

OpenStudy (anonymous):

:)

OpenStudy (kc_kennylau):

could u help me? :O

OpenStudy (anonymous):

sry i don't know math that complecated sry! :(

OpenStudy (anonymous):

sec

OpenStudy (anonymous):

i'll try

OpenStudy (anonymous):

sorry no matchs

OpenStudy (kc_kennylau):

lolz

OpenStudy (anonymous):

no no matchs i looked it up online

OpenStudy (anonymous):

i am trying it. gimme a few sec

OpenStudy (anonymous):

this is a creative solution contributed by trollswillrule: \[(1-\lceil \sin4\theta \rceil)(\csc \theta +\sec \theta) \]=f(x)=r

OpenStudy (anonymous):

for -45 degrees to 315 degrees

OpenStudy (anonymous):

this is to see that 1-the rounded up of sin4theta can become 0 and 1 in each case.

OpenStudy (anonymous):

just a clarification, if ceiling function of -0.1 is what? is it -1 or 0?

OpenStudy (kc_kennylau):

0

OpenStudy (anonymous):

Discrete Fourier transform is the best i can think of right now, i dunt think there are easier alternatives anymore...